### v1

@@ Abstract
In 2017, Bertsimas and Dunn proposed the first optimal decision tree (ODT) algorithm using general-purpose solvers, sparking significant interest in subsequent studies. However, existing combinatorial methods—primarily based on branch-and-bound (BnB) method—focus on much simpler variants, such as trees with binary feature data or axis-parallel hyperplanes. Although the first combinatorial algorithm appeared in 1989, little subsequent work has directly built upon it.Researchers have proposed various algorithms with different data structures and frameworks, but few adopt a unified framework comparable to general-purpose solvers. Many so-called optimal algorithms rely on ad-hoc processes and bounding strategies, with complex pseudo-code reducing accessibility and ambiguously stated claims hindering independent validation. As a result, prior BnB algorithms often lack the rigor to prove optimality, and algorithms claiming to solve the same problem may yield inconsistent—or even incorrect—solutions. Ultimately, optimal algorithms demand rigorous scrutiny: they must provably guarantee the optimal solution for a given search space and objective.
@@
We argue that the root cause lies in the absence of a formal framework for designing optimal algorithms using combinatorial methods and standardized criteria for verifying correctness. Any algorithm claiming optimality must be formally proven. To address this, in the first paper (Part I) of this two-part series, we examined the ambiguity and informality in studies of ODT problems and discussed the consequences of relying on informal formalisms in optimal algorithm design. We then presented criteria that an optimal algorithm should satisfy to ensure a verifiable and transparent design process.
@@
Beyond identifying limitations, our primary goal—and the central contribution of this work—is to develop a foundational theory for solving ODT problems. In Part I, we introduce four novel definitions of the ODT problems: three for size-constrained trees and one for depth-constrained trees. These definitions are stated unambiguously through executable recursive programs, satisfying all criteria we propose for a formal specification. In this sense, they resemble the “standard form” used in the study of general-purpose solvers.
@@
Grounded in algebraic programming theory—a relational formalism for deriving correct-by-construction algorithms from specifications—we can not only establish the existence or nonexistence of dynamic programming solutions but also derive them constructively whenever they exist. Consequently, the four generic problem definitions yield four novel optimal algorithms for ODT problems with arbitrary splitting rules satisfies the axioms and objective functions of a given form. These algorithms encompass the known depth-constrained, axis-parallel ODT algorithm as special case, while providing a unified, efficient, and elegant solution for the general ODT problem.
In Part II, we present the first optimal hypersurface decision tree algorithm and provide comprehensive experiments against axis-parallel decision tree algorithms, including heuristic CART and state-of-the-art optimal methods. The results demonstrate the significant potential of decision trees with flexible splitting rules. Moreover, our framework is readily extendable to support algorithms for constructing even more flexible decision trees, including those with mixed splitting rules.
@@ 1 Introduction
However, with the rapid development of ML research, the greedy top-down optimization algorithms are no longer as effective compared with state-of-the-art ML algorithms. To improve the accuracy of decision tree models, two common approaches are typically adopted. The first is to find the globally optimal decision tree (ODT) for the given data. Alternatively, instead of restricting models to axis-parallel hyperplanes, more complex splitting rules can be applied. For example, a generalization of the classical decision tree problem is the hyperplane (or oblique) decision tree, which employs hyperplane decision boundaries to potentially simplify complex boundary structures. However, the optimization problems in both approaches are computationally challenging. For the optimal axis-parallel decision tree (AODT) problem, Laurent and Rivest (1976) showed that determining whether there exists an axis-parallel decision tree with cost less than or equal to $w$ for a given finite dataset and rule set is NP-complete. Consequently, finding an optimal axis-parallel decision tree for a given dataset is NP-hard. Similarly, for hyperplane decision trees, even the greedy top-down optimization approach is NP-hard, as the number of possible hyperplane splits for a dataset of size $N$ in dimension $D$ is $O\left(N^{D}\right)$, which grows exponentially with $D$.
Nevertheless, the stopping criteria of the CART algorithm recursively split a leaf until one of the following conditions is met: (1) the maximum tree depth is reached, (2) the maximum number of leaves is attained, (3) the minimum number of observations per leaf is violated, or (4) the minimum impurity reduction threshold is satisfied. Upon closer examination, the optimization problem that the CART algorithm attempts to solve differs from the one defined by Laurent and Rivest (1976), as the latter does not impose the constraints specified in CART. This raises an important question: is the optimization problem solved by CART also NP-hard? Interestingly, Ordyniak and Szeider (2021) showed that the decision tree problem with fixed size and depth is fixed-parameter tractable. In other words, if the size or depth of the tree is fixed, there exists a polynomial-time algorithm for solving the problem.
While the result of Ordyniak and Szeider (2021) leaves certain aspects underspecified—for example, some key definitions regarding the problem formulation—it nevertheless offers a valuable perspective on the tractability of decision tree optimization. As a result, the fixed-parameter tractability of ODT problems has led to a significant growth of research interest in recent years. One major line of work focuses on formulating the ODT problem as a mixed-integer programming (MIP) problem, which can then be solved using general-purpose optimization solvers. Following the first MIP formulation proposed by Bertsimas and Dunn (2017), a surge of subsequent studies has investigated this approach (Boutilier et al., 2023; Verwer and Zhang, 2017, 2019; Boutilier et al., 2022; Zhu et al., 2020) leveraging general-purpose MIP solvers. Notably, Bertsimas and Dunn (2017)’s algorithm can not only solve the AODT problem but also the optimal hyperplane decision tree problem.
An alternative line of research, which we refer to as combinatorial methods, seeks to employ approaches such as dynamic programming (DP), greedy strategies, or branch-and-bound (BnB) techniques to exhaustively explore the solution space while eliminating non-optimal solutions through bounding techniques. Unlike the advances made in the study of general-purpose solvers, where many practical solutions have been developed for both the AODT problem and the hyperplane decision tree problem. Although the first combinatorial algorithm for decision tree optimization was introduced as early as 1989 (Cox Jr et al., 1989), most subsequent studies have focused on simpler settings, such as the ODT problem over binary feature data. Only a few studies, such as Brița et al. (2025); Mazumder et al. (2022), have addressed the AODT problem, whereas algorithms for the hyperplane optimal decision tree problem remain unexplored.
Unlike the work of Bertsimas and Dunn (2017) on the ODT problem, which has inspired a substantial body of subsequent research, many later studies on combinatorial algorithms adopt distinct problem definitions. For instance, the approaches of Demirović et al. (2022); Verwer and Zhang (2019); Nijssen and Fromont (2007); Aglin et al. (2020); Nijssen and Fromont (2010) construct optimal decision trees by constraining the tree depth. In contrast, Zhang et al. (2023); Lin et al. (2020); Hu et al. (2019) propose algorithms that constrain the number of leaves, referring to the resulting models as sparse decision trees.
Additionally, the combinatorial methods developed to solve the ODT problem are highly diverse. For example, Nijssen and Fromont (2007); Aglin et al. (2021, 2020); Nijssen and Fromont (2010) characterize decision trees over binary feature data using the concept of itemsets from data, Demirović et al. (2022) represent decision trees via feature vectors, whereas Lin et al. (2020); Hu et al. (2019) define trees in terms of two sets of leaves: those that can be further split and those that are fixed. Unfortunately, the advantages and limitations of these data structures remain insufficiently studied. If these distinct representations of decision trees prove to be indispensable—playing central roles in their respective algorithms and serving different purposes in tree construction—their strengths and weaknesses warrant careful analysis.
At the same time, while using general-purpose MIP solvers for the ODT problem is appealing and can be efficient in many cases, the overall computational complexity remains unpredictable. More importantly, MIP-based approaches provide limited insight into the combinatorial and geometric structure of the problem. Their performance largely depends on advances in general-purpose solvers and, potentially, more sophisticated MIP formulations. To advance theoretical understanding and develop practical algorithms, it is therefore essential to study combinatorial methods, which provide a more direct view of the underlying geometry and algorithmic structure.
In this study, we focus on combinatorial methods for solving the ODT problem, following the axiomatic definition of decision trees established in our previous work He and Little (2025b), which we refer to as proper decision trees, we provide a detailed analysis of the algorithmic, combinatorial, and geometric properties of the ODT problems. Our generic solution not only offers several new perspectives on decision tree optimization but also provides a rigorous framework for future analysis. Building on the algorithmic and geometric results discussed in this paper, another key contribution of this work is demonstrating that the hypersurface decision tree (HODT) model satisfies the axioms we proposed and can therefore be solved using the four ODT algorithms introduced here. Moreover, our algorithms are readily applicable to more flexible decision tree models that incorporate mixed splitting rules. In Part II, we will present the detailed construction of the first optimal hypersurface decision tree (HODT) algorithms, highlighting their practical implementation and applicability.
We believe this research lays a theoretical foundation for the study of ODT problems. A summary of our contributions is as follows:
Achievement in Part I
@@
  • Critique and proposed solutions: We critically examine current approaches to designing algorithms and defining problems in the study of optimal decision trees, highlighting the consequences of relying on informal problem formulations paired with ad-hoc algorithmic solutions—a practice common in combinatorial studies of ODT problems, particularly those employing BnB methods. Beyond this critique, the central distinction of our research lies in philosophy: we argue that deriving truly optimal algorithms demands the highest levels of rigor and clarity. Accordingly, we propose criteria for formally defining problems in the study of optimal algorithms, along with the key characteristics that a good optimal algorithm should possess. These discussions provide a foundational step for advancing research on ODT problems using combinatorial methods.
  • Formal combinatorial definitions: We provide four formal combinatorial definitions for the ODT problem—covering both size- and depth-constrained trees—ensuring unambiguous communication for future research. All definitions are directly executable, allowing verification that the generated trees align with expectations. To our knowledge, no prior studies on using BnB algorithms for ODT have provided a formal problem definition with this level of rigor.
@@
  • Generic optimal decision tree algorithms: We present four novel and generic algorithms for solving ODT problems with arbitrary splitting rules that satisfy our axioms, capable of optimizing arbitrary objective functions of a specified form. All algorithms are constructively derived in a formal manner. Specifically, we provide three algorithms for size-constrained trees and one for depth-constrained trees, formally proving and extending results from previous studies on axis-parallel ODT problems Brița et al. (2025).
@@
  • Filter and thinning fusion: We propose two fusion theorems—filtering and thinning—to enhance the flexibility and practical applicability of ODT algorithms while maintaining formal rigor. The filtering fusion theorem allows additional constraints, such as depth and leaf-size constraints, to be incorporated into ODT algorithms defined over size-constrained trees. Similarly, tree-size constraints can be applied to ODT algorithms defined over depth-constrained trees. The thinning fusion theorem integrates well-known bounding techniques into the framework. Together, these theorems extend the practical utility of the algorithms without compromising optimality.
@@
  • Proving the properness of hypersurface decision trees: By generalizing Xi’s 0-1 loss linear/hypersurface classification theorem from linear models to decision tree models, and combining it with the geometric characterization of polynomial hypersurfaces as combinations of data points, we demonstrate that hypersurface decision trees are proper decision trees and can therefore be solved exactly using our proposed algorithms.
@@
  • Splitting rule generators: Leveraging the sophisticated combination generator introduced by He and Little (2025a), we develop efficient algorithms for generating both hypersurface splitting rules and mixed-splitting rules for a given dataset. The hypersurface splitting rule generator provides the foundation for constructing efficient hypersurface decision tree algorithms developed in Part II, while the mixed-splitting rule generator establish a foundation for future work on constructing optimal ODT algorithms with mixed splitting rules.
@@
  • Complexity analysis for optimal decision tree problems: Most studies on ODT algorithms either do not analyze worst-case complexity or analyze it using incorrect combinatorial characterizations. Our analysis provides an upper bound for size-constrained decision trees, which also implies trivial lower bounds for depth-constrained axis-parallel decision tree problem. Moreover, this lower bound contradicts the function for counting the combinatorial complexity of the problem in Mazumder et al. (2022)’s source code, which may indicate a potential flaw regarding the exhaustiveness of their algorithm.
  • Challenge to existing algorithms: Because our algorithms are derived from a formal axiomatic definition of decision trees, they challenge prior studies that applied a similar algorithmic process to ODT problems with binary feature data (Demirović et al., 2022; Hu et al., 2019; Lin et al., 2020; Zhang et al., 2023; Aglin et al., 2021, 2020). These studies do not satisfy the axioms we propose, which are essential for justifying the correctness of a critical component, yet this component remain unspecified or ambiguously described in their constructions. It is therefore unclear whether their ambiguous algorithmic processes can support the claimed optimality of these algorithms, as none formally specified their problem or proved correctness. Our results call into question the validity of these claims.
Achievement Part II
  • Crossed-hyperplane properties: We identify the crossed-hyperplane properties, which significantly reduce the size of the search space for the HODT problem.
  • Incremental algorithm for ancestry matrix generation: Leveraging the prefix-closed property of crossed hyperplanes, we introduce an incremental algorithm for efficiently generating the ancestry relation matrix, which is crucial for solving size-constrained decision tree problems. This algorithm can be fully vectorized and massively parallelized.
  • Application of size-constrained ODT algorithms to HODT problem: We apply the two rigorously proven algorithms from Part I to solve the size-constrained HODT problem, encompassing classical axis-parallel hyperplanes, general hyperplanes, and hypersurfaces. Both algorithms have polynomial complexity (for fix tree size and data dimension), are easily parallelizable, and amenable to speed-up techniques.
  • Experiments on synthetic data: Using the two heuristic methods we proposed, we conduct comprehensive experiments on hyperplane decision algorithms with synthetic datasets, where the ground truth is generated by hyperplane decision trees (i.e., polygonal regions). The results show that hyperplane decision trees recover the ground truth more accurately than both heuristic methods and optimal axis-parallel tree algorithms under various noise conditions. Notably, our algorithms can even achieve solutions that are more accurate than the ground truth on noisy training data, refuting the notion that optimal algorithms necessarily overfit.
  • Benchmarking on real-world datasets: We perform a comprehensive analysis on 30 real-world datasetss, comparing hypersurface decision tree algorithms with CART and Brița et al. (2025)’s AODT algorithm (ConTree algorithm). The results demonstrate that hyperplane decision trees achieve superior generalization accuracy compared with axis-parallel decision trees when tree size is properly controlled.
@@ 2.1 Informality and consequences
Data back to the dawn of 20th century, after Gödel (1931) announced his well-known incompleteness theorem, there is a problem determining a formal definition of “effectively calculable.” To address this, three foundational models of computation were proposed: lambda calculus, invented by (Church, 1936), general recursive functions, proposed by Gödel (Gandy, 1988), and Turing machine, by Turing et al. (1936). Church give his definition of $\lambda$-definability and claim all calculation that is $\lambda$-definable are effective calculable. However, his argument was unsatisfactory to Gödel. As (Wadler, 2015) noted:“ Church merely presented the definition of $\lambda$-definability and baldly claimed it corresponded to effective calculability.”
@@
Unfortunately, a common trend in the studies of optimal algorithms using combinatorial methods are to focus on proving bounding techniques without establishing the correctness of the main algorithm. However, establishing the optimality of an algorithm requires more than just validating a bounding technique. In essence, much of this work amounts to the design of “optimal bounding techniques” rather than the development of truly “optimal algorithms” for the decision tree problem. These are fundamentally different research directions, as a bounding technique is meaningful only when it supports a provably correct algorithm in the first place. We believe this inclination stems from a programming philosophy that prioritizes efficiency as the primary, and sometimes sole, objective. This tendency is further compounded by the lack of widely accepted criteria for defining an optimal algorithm.
@@
For instance, studies of ODT algorithms using combinatorial methods often leave the problem undefined [^1] We claim that these studies leave the problem undefined because they commonly do not specify the solution space over which they are searching for an optimal decision tree. In some cases, the space is not even ambiguously defined, as in specification (1), making it unclear which ODT problem they are actually attempting to solve. (Lin et al., 2020; Zhang et al., 2023; van der Linden et al., 2023; Demirović et al., 2022; Hu et al., 2019) or specify their problem ambiguously, as in (Mazumder et al., 2022; Brița et al., 2025), where the problem is expressed in a form similar to
@@
Although (1) appears intuitively obvious, it suffers from several major flaws. First, the search space $\mathcal{S}$ usually denoted by a symbol and defined through text description. For example, Brița et al. (2025) wrote:“ $\mathcal{T}\left(\mathcal{D},d\right)$ describe the set of all decision trees for the datasets $\mathcal{D}$ with a maximum depth of $d$ .” Adopting such a textual description for $\mathcal{T}\left(\mathcal{D},d\right)$ is inherently ambiguous and dangerous. Even though it appears simple, different researchers may interpret the search space differently due to variations in the constraints imposed on the trees, and some may even adopt entirely different definitions of the decision tree. In contrast, MIP specifications force researchers to define precisely what they mean by a decision tree by specifying the constraints first, enabling an unambiguous definition.
As a result, studies that adopt such informal problem definitions cannot verify the most important step in proving the optimality of an algorithm—showing that the algorithm provably searches all possible solutions in $\mathcal{S}$ ($\mathcal{T}\left(\mathcal{D},d\right)$ in Brița et al. (2025)), because $\mathcal{S}$ is not formally defined. Consequently, research on BnB algorithms often focuses solely on proving the correctness of the bounding techniques and then claims the algorithm is “optimal.”
Additionally, the ambiguous problem definition inevitably leads to inconsistencies in the study of optimal algorithms. For example, although Mazumder et al. (2022) and Brița et al. (2025) claim to solve the same problem, we found that their algorithms return different trees with significantly different objective values on the same datasets—differences too large to be attributed to numerical issues. This suggest that either Mazumder et al. (2022)’s algorithm is non-optimal or that the two studies are solving different ODT problems. Evidence favors the first explanation: a function in Mazumder et al. (2022)’s source code counts the number of possible trees in the search space, which contradicts the lower bound we provide in Subsection 5.3.

### v2

@@ Abstract
In 2017, Bertsimas and Dunn proposed the first optimal decision tree (ODT) algorithm using general-purpose solvers, sparking significant interest in subsequent studies. However, existing combinatorial methods—primarily based on branch-and-bound method—focus on much simpler variants, such as trees with binary feature data or axis-parallel hyperplanes. Although the first combinatorial algorithm appeared in 1989, little subsequent work has directly built upon it. Researchers have proposed various algorithms with different data structures and frameworks, but few adopt a unified framework comparable to general-purpose solvers. Many so-called optimal algorithms rely on ad-hoc algorithmic processes and bounding strategies, with complex pseudo-code reducing accessibility and ambiguously stated claims hindering independent validation. However, optimal algorithms require rigorous verification—they must provably guarantee optimal solutions for a given search space and objective. Previous BnB algorithms often fail to meet this level of rigor, rendering claims of fully reliable optimality, at best, premature.
@@
We argue that the root cause lies in the absence of a formal framework for designing optimal algorithms using combinatorial methods and standardized criteria for verifying correctness. Any algorithm claiming optimality must be formally proven. To address this, in the first paper (part I) of this series of two, we examined the ambiguity and informality in studies of ODT problems and discussed the consequences of relying on informal formalisms in optimal algorithm design. We then presented criteria that an optimal algorithm should satisfy to ensure a verifiable and transparent design process.
@@
Beyond identifying limitations, our primary goal—and the central contribution of this work—is to develop a foundational theory for solving ODT problems. We aim to establish a comprehensive algorithmic framework for all optimal decision tree problems studied in machine learning, encompassing axis-parallel, hyperplane, hypersurface decision trees, and those defined over binary feature data. In Part I, we introduce five rigorous formulations of ODT problems: three for size-constrained trees and one for depth-constrained trees, and one for trees over binary feature data. These definitions are stated unambiguously through executable recursive programs, satisfying all criteria we propose for a formal specification. In this sense, they resemble the “standard form” used in the study of general-purpose solvers.
@@
Grounded in algebraic programming theory—a relational formalism for deriving correct-by-construction algorithms from specifications—we can not only establish the existence or nonexistence of dynamic programming solutions but also derive them constructively whenever they exist. Consequently, the five generic problem definitions yield five optimal algorithms for ODT problems with arbitrary splitting rules satisfies the axioms and objective functions of a given form. These algorithms encompass the known ODT algorithms as special case, while providing a unified, efficient, and elegant solution for the general ODT problem.
Keywords: decision tree, dynamic programming, global optimal algorithm
@@ 1 Introduction
However, with the rapid development of ML research, the greedy top-down optimization algorithms are no longer as effective compared with state-of-the-art ML algorithms. To improve the accuracy of decision tree models, two common approaches are typically adopted. The first is to find the globally optimal decision tree (ODT) for the given data. Alternatively, instead of restricting models to axis-parallel hyperplanes, more complex splitting rules can be applied. For example, a generalization of the classical decision tree problem is the hyperplane (or oblique) decision tree, which employs hyperplane decision boundaries to potentially simplify complex boundary structures.
However, the decision tree optimization is notoriously difficult. Laurent and Rivest (1976) showed that finding the smallest decision tree (defined by the path lengths) is NP-complete. Consequently, finding an optimal axis-parallel decision tree for a given dataset is NP-hard. Similarly, for hyperplane decision trees, even the greedy top-down optimization approach is NP-hard, as the number of possible hyperplane splits for a dataset of size $N$ in dimension $D$ is $O\left(N^{D}\right)$, which grows exponentially with $D$.
However, instead of addressing the general decision tree problem that seeks the smallest tree, successful algorithms such as CART and C4.5 were not designed with this objective. Rather, their stopping criteria explicitly constrain either the tree size (i.e., the number of branch nodes) or the tree depth. Interestingly, Ordyniak and Szeider (2021) demonstrated that the decision tree problem with fixed size or depth is fixed-parameter tractable. In other words, when either parameter is fixed, the problem can be solved in polynomial time.
In this first paper of a two-part series, we present a framework for constructing efficient algorithms—polynomial in the worst case and trivially parallelizable—for obtaining globally optimal solutions to decision tree optimization problems. Our exposition offers solutions to nearly all ODT problems encountered in practice, including classical axis-parallel hyperplane, general hyperplane, polynomial hypersurface, and decision trees over binary feature data. In Part II (He, 2025a), we will detail the construction of the first optimal hypersurface decision tree algorithms, emphasizing their practical implementation and applicability.
As our goal is to provide a comprehensive algorithmic framework for all decision tree problems relevant in practice, this paper is written in a pedagogical style. Nonetheless, it includes several novel contributions. A summary of our main contributions in Part I is as follows:
@@
  • Critique and proposed solutions: In Section 2.1, We critically examine current approaches to designing algorithms and defining problems in the study of optimal decision trees, highlighting the consequences of relying on informal problem formulations paired with ad-hoc algorithmic solutions—a practice common in combinatorial studies of ODT problems, particularly those employing BnB methods. Beyond this critique, the central distinction of our research lies in philosophy: we argue that deriving truly optimal algorithms demands the highest levels of rigor and clarity. Accordingly, we propose criteria for formally defining problems in the study of optimal algorithms, along with the key characteristics that a good optimal algorithm should possess. These discussions provide a foundational step for advancing research on ODT problems using combinatorial methods.
  • Formal combinatorial definitions: Following the proper decision tree framework examined in He and Little (2025b), which defines decision trees through recursive datatypes and establishes four axioms, In Subsection 4.2, we established four formal combinatorial definitions for the ODT problem—covering both size- and depth-constrained trees. Importantly, we rigorously defined the search space and the objective function of the ODT problems as directly executable programs. This is vitally important because it allows researchers to inspect the search space by running the program and verifying the resulting decision trees. Such transparency not only validates our solution but also fosters unambiguous communication. To our knowledge, no prior work on BnB algorithms for ODT problems has provided a problem definition with this level of rigor, combining axiomatic definitions of decision trees with programmatic definitions of the search space and objective function.
@@
  • Generic optimal decision tree algorithms: In Subsection 4.3, We present four generic algorithms for solving ODT problems with arbitrary splitting rules over proper decision trees, capable of optimizing arbitrary objective functions of a specified form. One algorithm was previously introduced in our earlier work He and Little (2025b), while the remaining three are new contributions in this exposition. All algorithms are derived constructively and presented in a formal manner. Specifically, we provide three algorithms for size-constrained trees and one for depth-constrained trees. Notably, our depth-constrained ODT algorithm not only offers a formal proof of the results in Brita et al. (2025) for the axis-parallel ODT problem but also generalizes their findings to broader settings.
  • Extension to decision tree over binary feature data: He and Little (2025b) showed that the ODT-BF problem is a non-proper decision tree problem, requiring a distinct formulation of its search space. Based on this formulation, we derived a provably optimal algorithm for solving it. To the best of our knowledge, this is the first optimal algorithm that guarantees exhaustive exploration of the ODT-BF search space, thereby establishing a solid foundation for addressing this problem.
  • Filter and thinning fusion: In Subsection 4.5, We propose two fusion theorems—
@@
    filtering and thinning—to enhance the flexibility and practical applicability of ODT algorithms while maintaining formal rigor. The filtering fusion theorem allows additional constraints, such as depth and leaf-size constraints, to be incorporated into ODT algorithms defined over size-constrained trees. Similarly, tree-size constraints can be applied to ODT algorithms defined over depth-constrained trees. The thinning fusion theorem integrates well-known bounding techniques into the framework. Together, these theorems extend the practical utility of the algorithms without compromising optimality.
@@
  • Proving the properness of hypersurface decision trees: In Section 5, by generalizing He and Little (2023)’s 0-1 loss linear/hypersurface classification theorem from linear models to decision tree models, and combining it with the geometric characterization of polynomial hypersurfaces as combinations of data points, we demonstrate that hypersurface decision trees are proper decision trees and can therefore be solved exactly using our proposed algorithms.
@@
  • Splitting rule generators: In Subsection 5.2, leveraging the sophisticated combination generator introduced by He and Little (2025a), we develop efficient algorithms for generating both hypersurface splitting rules and mixed-splitting rules for a given dataset. The hypersurface splitting rule generator provides the foundation for constructing efficient hypersurface decision tree algorithms developed in Part II, while the mixed-splitting rule generator establish a foundation for future work on constructing optimal ODT algorithms with mixed splitting rules.
@@
  • Complexity analysis for optimal decision tree problems: In Subsection 5.3, Most studies on ODT algorithms either do not analyze worst-case complexity or analyze it using incorrect combinatorial characterizations. Our analysis provides an upper bound for size-constrained decision trees, which also implies trivial lower bounds for depth-constrained axis-parallel decision tree problem. Our lower bound analysis reveals a discrepancy with the function used to count the combinatorial complexity in the source code of Mazumder et al. (2022). This observation suggests a possible limitation in the exhaustiveness of their algorithm, which could warrant further investigation to refine the approach.
@@ 2.1 Informality and consequences
Data back to the dawn of 20th century, after Gödel (1931) announced his well-known incompleteness theorem, there is a problem determining a formal definition of “effectively calculable.” To address this, three foundational models of computation were proposed: lambda calculus, invented by (Church, 1936), general recursive functions, proposed by Gödel (Gandy, 1988), and Turing machine, by Turing et al. (1936). Church give his definition of $\lambda$-definability and claim all calculation that is $\lambda$-definable are effective calculable. However, his argument was unsatisfactory to Gödel. As Wadler (2015) noted:“ Church merely presented the definition of $\lambda$-definability and baldly claimed it corresponded to effective calculability.”
@@
Unfortunately, a common trend in the studies of optimal algorithms using combinatorial methods are to focus on proving bounding techniques without establishing the correctness of the main algorithm. However, establishing the optimality of an algorithm requires more than just validating a bounding technique. In essence, much of these works amounts to the design of “optimal bounding techniques” rather than the development of truly “optimal algorithms” for the decision tree problem. These are fundamentally different research directions, as a bounding technique is meaningful only when it supports a provably correct algorithm in the first place. We believe this inclination stems from a programming philosophy that prioritizes efficiency as the primary, and sometimes sole, objective. This tendency is further compounded by the lack of widely accepted criteria for defining an optimal algorithm.
@@
For instance, studies of ODT algorithms using combinatorial methods often leave the problem undefined [^1] We claim that these studies leave the problem undefined because they commonly do not specify the solution space over which they are searching for an optimal decision tree. In some cases, the space is not even ambiguously defined, as in specification (1), making it unclear which ODT problem they are actually attempting to solve. (Lin et al., 2020; Zhang et al., 2023; van der Linden et al., 2023; Demirovic et al., 2022; Hu et al., 2019) or specify their problem ambiguously, as in (Mazumder et al., 2022; Brita et al., 2025), where the problem is expressed in a form similar to
@@
Although (1) appears intuitively obvious, it suffers from several major flaws. First, the search space $\mathcal{S}$ usually denoted by a symbol and defined through text description. For example, Brita et al. (2025) wrote:“ $\mathcal{T}\left(\mathcal{D},d\right)$ describe the set of all decision trees for the datasets $\mathcal{D}$ with a maximum depth of $d$ .” Adopting such a textual description for $\mathcal{T}\left(\mathcal{D},d\right)$ is inherently ambiguous and dangerous. Even though it appears simple, different researchers may interpret the search space differently due to variations in the constraints imposed on the trees, and some may even adopt entirely different definitions of the decision tree. In contrast, MIP specifications force researchers to define precisely what they mean by a decision tree by specifying the constraints first, enabling an unambiguous definition.
As a result, studies that adopt such informal problem definitions cannot verify the most important step in proving the optimality of an algorithm—showing that the algorithm provably searches all possible solutions in $\mathcal{S}$ ($\mathcal{T}\left(\mathcal{D},d\right)$ in Brita et al. (2025)), because $\mathcal{S}$ is not formally defined. Consequently, research on BnB algorithms often focuses solely on proving the correctness of the bounding techniques and then claims the algorithm is “optimal.”
Additionally, the ambiguous problem definition inevitably leads to inconsistencies in the study of optimal algorithms. For example, although Mazumder et al. (2022) and Brita et al. (2025) claim to solve the same problem, we found that their algorithms return different trees with significantly different objective values on the same datasets—differences too large to be attributed to numerical issues. This suggest that either Mazumder et al. (2022)’s algorithm is non-optimal or that the two studies are solving different ODT problems. Evidence favors the first explanation: a function in Mazumder et al. (2022)’s source code counts the number of possible trees in the search space, which contradicts the lower bound we provide in Subsection 5.3.